OpenStudy (anonymous):

@shamim @mialovesyou03

OpenStudy (anonymous):

(6-3i)/(2+1)

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

do you mean divide or a fraction?

OpenStudy (anonymous):

fraction

OpenStudy (anonymous):

Simplify each expression. Write your final answer in a+bi form. Show your work

OpenStudy (anonymous):

wait don't you mean 2+i in the question?

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

are you sure thats a 1? cause if it it doesn't make sense if it is, I only know how to solve it if thats an i

OpenStudy (anonymous):

its an i

OpenStudy (anonymous):

First off, the common conjugate is (2-i) so the next step is (6-3i)(2-i) / (2+i)(2-i) [6*2 + 6*(-i)-3i*2-3i*-i] + [2*-i+i*2+i*-i] then simplify then cut down numbers (12+(-6-6i)*i+3(-1)) / (4+(-2+2)i-(-1)) (12-12i-3) / (4+1) a=9/5 b=-12/5 HOPE YOU UNDERSTAND THAT :)

OpenStudy (anonymous):

thanks:)